# Why Johnny Signs with Sigstore: Examining Tooling as a Factor in Software Signing Adoption in the Sigstore Ecosystem

### v3

@@ 4.1.1 Rationale
The case study is a research method used to investigate a phenomenon within its real-life context  [61, 91, 18, 24]. We selected the case study method for our research because it allows for an in-depth exploratory examination of our chosen case, enabling a deeper understanding of the usability factors that influence the adoption of a software signing tool in practice, rather than a broad but superficial cross-analysis  [59]. In selecting Sigstore, we adhered to the criteria outlined by Yin [91] and Crowe et al. [18]. We considered the intrinsic uniqueness of the unit (high), access to data (adequate), and risks of participation (low).
Our research objectives explore the impact of Software Tooling as a factor in the practice of software signing from the viewpoint of practitioners, with a particular focus on how the usability of software signing tools influences their adoption. We opted for semi-structured interviews [66] to effectively examine various aspects of tooling, including usability challenges, perceived benefits, and comparative experiences with other signing tools. This approach allows us to gather descriptive, open-ended, opinionated, and detailed responses, providing deeper insights into how usability considerations shape adoption decisions.
4.1.2 Instrument Design & Development
To develop our interview instrument, we relied on the usability framework (Figure 1) and various signing literature. We outline how these resources informed our instrument design:
@@ 4.1.2 Instrument Design & Development
    Grey Literature: To derive our questions, we primarily examined Sigstore blogs [74], and Sigstore usage reports [8, 86, 31, 38, 76, 75, 77, 15]. These sources included reports on how various entities implemented Sigstore in their engineering processes, news about major updates, and other related Sigstore developments.
@@
    Background & Literature Review (section 2): We employed a snowball approach to examine works from top security conferences on software signing, the design of software signing primitives [53, 52], and empirical studies on software signing (all cited in (section 2)). Our literature review informed us of existing knowledge gaps and provided insights into the design of modern software signing tools. This helped us avoid duplicating questions already addressed in the literature.
  3. 3.
    We designed our questions to align with our formative usability framework (section 3). We developed these questions to focus on identifying the factors responsible for the adoption of Sigstore in practice. Categories D3 and D4 focused on participants’ experiences with their current tools, aligning with the framework’s emphasis on assessing present tooling experiences. Meanwhile, categories D1, D2, and D5 captured participants’ experiences with previous software signing tools, offering insights into their transitions and comparisons over time.
Next, we refined our interview protocols through a series of practice and pilot interviews. We first conducted two practice interviews with the secondary authors of this work, followed by pilot interviews with the first two interview subjects.
Our full interview instrument consisted of three main topics (B-D). As suggested by the table, topics B and C do not fit into the same analysis framework as topic D. Thus, following the guidance of Voils et al. [88], we analyzed those topics in another work [37]. In this work, we focus on the data for Topic D. Topic D consisted of five main question categories, conducted in a semi-structured manner. A summary of the questions in our interview protocol is shown in Table 1.
Table 1: Summary of interview protocol. Topic B and C are analyzed in [37]. We analyze only topic D in this paper. The decision to analyze these parts of the protocol separately was made in accordance with best practices. We present our full protocol in  Appendix A
@@ D. Signing tool adoption (5)
4.2 Data Collection
Recruitment: Our study investigates signing tools using a case study of the Sigstore ecosystem. As such, our target participant pool is centered on Sigstore users. Furthermore, our research questions, as reflected in our interview instrument, aim to understand the reasons prior to adoption that teams choose software tools (Sigstore), how Sigstore is implemented and the challenges and benefits of Sigstore.
@@ 4.2 Data Collection
To recruit this target population, we employed a non-probability purposive and snowball sampling approach  [7]. We leveraged the personal network of one of the authors to send an initial invitation to members of the 2023 KubeCon (hosted by the Linux Foundation) organizing committee, who were also part of the In-toto [83, 32] Steering Committee (ITSC). This initial invitation yielded six (6) participants. We then expanded our recruitment through snowball sampling, relying on recommendations from initial participants (five(5)) and authors contacts (seven) to recruit an additional twelve(12) participants. We interviewed 18 subjects, but only responses from 13 participants fit our analysis framework [^1] We interviewed 18 participants but ultimately excluded data from five for this study. Two of them used software signing tools other than Sigstore, and three gave limited information about their software signing tool..
Participant Demographics:
Table 2: Subject Demographics. For anonymity, we used generic job roles, not specific titles. “Senior management” refers to senior managers, directors, and executives; “Technical leader” to senior, lead, partner, and principal engineers; and “Engineer” and “Manager” to more junior staff. We highlight type of software subject’s immediate team produces. * refers to cases where subject’s team built other types of software in addition to the main stated. O refers to cases where subject belonged to multiple teams with other products to the one stated. We also highlight which participants were recruited from the In-toto Steering committee (I) and which were from snowball sampling (S).
ID Role Experience Software Type Rec.
P1 Research leader 5 years Internal POC software I
P2 Senior mgmt. 15 years SAAS security tool*O S
P3 Senior mgmt. 13 years SAAS security tool*O I
P4 Technical leader 20 years Open-source tooling S
P5 Engineer 2 years Internal security tooling S
P6 Technical leader 27 years Internal security tooling*O S
P7 Manager 6 years Security tooling*O S
P8 Engineer 2.5 years SAAS security* I
P9 Technical leader 4 years Consultancy S
P10 Research leader 13 years POC security Software* S
P11 Senior mgmt. 15 years Internal security tooling I
P12 Senior mgmt. 15 years SAAS S
P13 Manager 11 years Security tooling S
We summarize our interview participants’ demographic information in Table 2. Our interview sampling approach yielded 13 experienced security practitioners who indicated that they are responsible for initiating or implementing their organization’s security controls and (or) have compliance-related responsibilities. This involvement gives them the relevant context to assess their organization’s strategies for adopting software signing tools.   Interviews: We conducted our interviews over Zoom. Each interview lasted a median of 50 minutes. The lead author conducted these interviews. We offered each subject a \$100 gift card as an incentive, in recognition of their experience level.
4.3 Data Analysis
@@
To analyze our interviews, we employed a thematic analysis approach [9] alongside our modified usability framework (section 3, Figure 1) within an interpretivist paradigm [4].
An interpretivist approach allows each participant’s response to be treated as a unique case, considering its specific context and surrounding variables. This is particularly valuable in our study, as participants belong to different teams with diverse implementation contexts and varying uses of Sigstore.
@@
Our analysis began with a data familiarization stage [80] and proceeded through three phases, carried out by two analysts  [^2] One of the analysts is the lead author.. In the first two phases, we inductively memoed and coded our transcripts, developing our initial codebooks without referencing our usability framework [21]. In the final phase, we derived themes from our inductively generated codes and then applied our usability framework to further refine the extracted themes. We describe these steps in detail next.
@@
Memoing & Codebook Creation: First, beginning with the initial 18 transcripts, the two analysts independently memoed a subset of six anonymized transcripts across three rounds of review and discussion, following the recommendations and techniques outlined by O’Connor & Joffe [56] and Campbell et al. [12]. This approach ensured the reliability of our analysis process while optimizing data resources. [^3] Although there is no consensus on the ideal proportion of data to use, O’Connor and Joffe suggest that selecting 10–25% is typical [56]. We randomly selected approximately 33% (six transcripts) to develop our codebook. This slightly exceeded O’Connor & Joffe’s recommended percentage to allow the analysts a better familiarization with the nature of our collected data, and analysis format.
After each round, emerging codes were discussed in meetings, resulting in a total of 506 coded memos.
@@
We then reviewed the coded memos and finalized an initial codebook with 36 code categories. To assess the reliability of the coding process, we combined the techniques outlined by Maxam & Davis [51] and Campbell et al. [12]. Specifically, we randomly selected an unanalyzed transcript, which we coded independently. Following Feng et al. ’s [20] recommendation, we used the percentage agreement measure to evaluate consistency, as both analysts had contributed to developing the codebook. We obtained an 89% agreement score and resolved all disagreements through discussions and code refinements. This high agreement indicates the stability of our coding scheme, allowing us to reliably singly code the remaining transcripts [12, 56].
@@
Codebook Revision: Following the reliability assessment of our codebook, and given the substantive agreement recorded, one of the analysts proceeded to code the remaining transcripts in two batches of six transcripts per round. In the first batch, six new code categories were introduced, which were subsequently reviewed by the other analyst. The same analytical process was applied to the final six transcripts, leading to the addition of four new code categories, which were subsequently reviewed by the other analyst. Notably, of the 10 codes added during this phase of codebook refinement, only three provided new insights into the data. By mutual agreement between both analysts, these were categorized as minor codes.
@@
To ensure our sample size and analysis was adequate, we calculated code saturation for each category in our codebook. Following the recommendations and methods of Guest et al. [26] and Maxam & Davis [51], we measured the cumulative number of new codes introduced in each interview. The results are presented in Figure 6. As shown, saturation was achieved after interview 11 , with no new codes emerging beyond this point. Each interview had a median of 15 unique codes (for topic D alone), as illustrated in Figure 6.
Thematic Analysis & Usability Framework Analysis: Next, following the method proposed by Braun & Clarke [9], we derived initial themes from our code categories. We focused on identifying themes related to the D part of our protocol, as highlighted earlier in Table 1. We then conducted another round of thematic analysis, incorporating the initial themes alongside our formative usability framework (section 3). In this phase, we mapped our initial themes onto the factors defined within our framework.
Survey Attempt: In addition to our interviews, we designed and conducted a survey to gather insights from members of the Sigstore community who might not have had the time for an interview. However, our survey received only a few responses, highlighting the challenges of conducting surveys within highly specialized communities [78]. We elaborate on our survey results and details in Appendix D.
4.4 Study Limitation & Threats to Validity
To discuss the limitations of our work, we follow Verdecchia et al.’s [87] recommendation to reflect on the greatest threats to our research. In addition, we discuss how our experiences and perspectives may have influenced our research.
Construct Validity: Cresswell’s usability framework was developed for health information technology. We adapted it to software signing tooling, but may have introduced flaws.
External Validity: The primary bias to our work is that we recruited from a population biased toward the use of Sigstore. We viewed this as necessary in order to recruit sufficient subjects, and feel that that result — a focused treatment of usability in Sigstore — is a valuable contribution to the literature. A secondary threat arises from the sample size of our work, which is due to the challenges of recruiting expert subjects. A final threat to the validity of our research is that it reflects the state of Sigstore at only one point in time. Our interviews were conducted between November 1, 2023, and February 9, 2024 and reflect the experiences of Sigstore users at that time.
Statement of Positionality: We acknowledge that one of our paper’s authors is a Sigstore tool contributor. This author was not involved in developing the research design nor the initial analysis of the results, but provided insight in later analysis.
5 Results
5.1 RQ1: Problems & Strengths of Sigstore
@@
Table 3: Summary of Practitioner-Reported Advantages of Using Sigstore
@@
Signing Workflow & Verification P1, P2, P7, P8, P10-13
@@
Use of Short-lived Keys & Certificate 3 –P2, P3, P11
@@
Integrability with SLSA build P12
Integrability with several container registries/technologies P9, P10
Integrability with several cloud-native applications P11
@@
Transparency logs increase security P5, P10
Evaluation of signing adoption using logs P8
@@
Free/Open-Source 2 – P7, P13
The first aim of our research work was to ascertain from the the experiences of our subjects the problem and strengths they have experienced while using Sigstore. We categorized each identified theme based on our usability framework factors: Technological Factors – situations where observed conditions were linked to the behavior of Sigstore technology itself, Human/Social Factors – participants’ observations related to gaps in support resources, previous experiences (with other tools) Organizational Factors – aspects tied to users’ organizations or their organizational capacities, and Macroenvironmental Factors – external influences affecting adoption and use.
5.1.1 Sigstore’s Strength
We present a summarized list of practitioner-reported advantages of using Sigstore in Table 3. First, we extracted the benefits our participants experienced while using Sigstore from their responses. These reflections capture their firsthand experiences with Sigstore.
We grouped their responses into eight categories: ease of use, identity management for signers, compatibility with emerging technologies, use of short-lived keys and certificates, implementation of a transparency log, encouragement of bundling signatures with attestations, free access to the public Sigstore instance, and reliability of the Sigstore service. When mapped to our modified usability framework, these factors fell under two categories: Technological and Macroenvironmental factors from our framework.
@@ 5.1.1 Sigstore’s Strength
Ease of use: Ease of Use was a predominant reason why many participants chose to use Sigstore, making it the most frequently cited advantage over other Software signing tools. Participants commonly highlighted the simplicity of following the Sigstore workflow for signing and verifying signatures. P7 (security tools, cloud security) states, “ I don’t really know of any other options that provide the same conveniences that Sigstore provides… It really is just one command to sign something and then one more command to verify it, and so much of the work is handled by Sigstore.” Other advantages associated with Ease of Use, as mentioned by practitioners, include the seamless integration of the Sigstore workflow into CI/CD automation using GitHub Actions and the reduced complexity of key management.
Identity management for Signers: Sigstore’s integration of OIDC identity management was a key strength reported by our participants. In traditional software signing, signer identity is often an ad-hoc feature that may or may not be required by the signing authority. In Sigstore, however, this feature—implemented via OIDC—is integrated with the Fulcio certificate authority as an identity management step before creating a signature. Participants greatly valued the elimination of the key generation and distribution phase, as it reduced the time spent distributing keys to key servers and searching for keys, streamlining the signing process. P12 (SAAS, SSC security tool) Highlights this point, “The other strength that it has is I can associate my identity with an OIDC identity as opposed to necessarily needing to generate a key and keep track of that key … that’s super useful because I could say, "Oh, this is signed with Mike’s public GitHub identity." So unless my GitHub identity has been compromised, that’s much better.”
Compatibility with Several New Technologies: A portion of our participants reported that Sigstore’s integration and compatibility with a wide range of modern technologies provided a significant advantage. The mentioned technologies included security tooling such as SLSA build attestations, as well as containerization technologies and cloud computing applications. In most cases, the resources required to set up connections to these technologies were minimal or nonexistent. P9 (consultancy, cloud OSS security) states, “a lot of clients are getting into signing their containers pretty trivial to do if you use something like cosign…Sigstore is the easy choice because it just works on any registry, and integrating it to Kubernetes is pretty easy ”
@@
Presence of a Transparency Log: The inclusion of a transparency log in the Sigstore ecosystem, serving as a tamper-resistant secondary record of changes to an artifact’s metadata, was seen as highly beneficial by participants. This feature provided an additional layer of trust, allowing users to verify signed artifacts with greater confidence. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P8 (SAAS, SSC security): “From a higher level, we’re going to use a transparency log to determine what was actually signed and what was written into that transparency log.”
Other technological factors noted by participants are the reliability of Sigstore’s public instance was a valuable aspect of Sigstore.
@@
Free/Open-Source:
Sigstore’s open-source nature was also highlighted by participants as a key advantage. They noted that the no-cost solution reduces both setup and maintenance barriers, particularly for users leveraging Sigstore’s public deployments. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P7 (security tools, cloud security): “I think it’s amazing that Sigstore is free and public”
5.1.2 Sigstore’s Weaknessses
We summarize the difficulties reported by participants in using Sigstore in  Table 4. Most of the weaknesses reported often belong to multi-factors. e.g., In the problems associated with Enterprise applicabiliirty of Sigstore, Some participants reported this as a consequence of regulations place on organizations which are both macroenvironmental (regulations) and Organizational.
Table 4: Reported Difficulties Encountered by Participants in Using Sigstore. We indicate the associated usability factor of each category of weakness using – T-technology, P-social/human, O-organizational, M-macroenvironmental factors.

### v4

@@ IV-A Rationale for Case Study
The case study is a research method used to investigate a phenomenon within its real-life context  [58, 59, 60, 61]. We selected the case study method for our research because it allows for an in-depth exploratory examination of our chosen case, enabling a deeper understanding of the usability factors that influence the adoption of a software signing tool in practice, rather than a broad but superficial cross-analysis [62]. In selecting Sigstore, we adhered to the criteria outlined by Yin [59] and Crowe et al. [60]. Specifically, we considered the intrinsic value and uniqueness of the case study unit (high), access to data (adequate), and risks of participation (low). In section VI we discuss how our findings may generalize.
Figure 2: Study Methodology. We developed our interview protocol by reviewing Sigstore grey literature and the academic literature. Finally, we analyzed our results using a thematic analysis informed by Cresswell’s formative usability evaluation framework (Figure 1).
IV-B Case Study Context – Sigstore
Following guidelines outlined by Runeson & Höst [58], we describe the importance of our selected case study context and the technical details of the system.
Use in Practice: Sigstore launched in 2021. It provides keyless, ephemeral keys and signer authentication mechanisms to simplify the signing process used by traditional tools. Within 13 months of its public release, Sigstore recorded 46 million artifact signatures [25]. It has significant adoption and support from open-source ecosystems (e.g., PYPI [26], Maven [24], NPM [63]) and many companies (e.g., Autodesk [31], Verizon [29], Yahoo[64]). Thus, Sigstore is a major next-generation software signing platform, making it a suitable context for examining usability in modern signing tools.
Components & Workflow: We detail Sigstore’s architecture and signing workflow, to familiarize the reader with this example of next-generation software signing tools, and to interpret participants’ statements referring to these components.
Sigstore’s primary components are [11]:
  • Cosign: CLI and Go library to create, verify, and bundle signatures and attestations.
  • OIDC Identity Provider: OpenID Connect services (e.g., GitHub, Google, Microsoft) that authenticate signers and issue short-lived identity tokens.
  • Fulcio (Certificate Authority): Issues ephemeral signing certificates after verifying the signer’s OIDC token.
  • Rekor (Transparency Log): Append-only ledger that records each signing event and certificate issuance.
  • Monitors and Verifiers: Services or tools that validate the signer’s identity, certificate status, and signature integrity.
Figure 3 shows the Sigstore software signing workflow. First, the signer authenticates with an OIDC provider to obtain a short-lived token. Next, Fulcio verifies this token and issues a one-time signing certificate. The signer then uses Cosign to apply that certificate to the software artifact. Finally, Rekor logs the certificate and signature, and verifiers retrieve entries from both Fulcio’s log and Rekor to confirm the signer’s identity and ensure the certificate and signature remain valid.
Figure 3: Sigstore Signing Workflow. The software author requests a certificate from a certificate authority, which confirms the signer’s identity through an identity provider. The signature and signer’s identity are recorded in a transparency log, which the verifier can monitor to confirm the validity of the signature upon downloading the signed package. See section -K for legible figure
Commonality with other Next-Gen Tools: Sigstore, like other next-generation signing tools such as OpenPubKey [37], AWS Signer [65], and SignServer [38], shares core functionality in signer identity management and automated key handling. However, unlike Sigstore, most of these tools do not include a built-in certificate authority (e.g., OpenPubKey) or a transparency log (e.g., OpenPubKey, SignServer). Instead, such services must be added via external integrations. While our focus on Sigstore allows for an in-depth case analysis, many of its design principles—such as identity-based authentication and ephemeral keys—are shared by its peers, suggesting potential generalizability across next-generation signing tools
IV-C Instrument Design & Development
Following guidance from Saldaña [66], we used semi-structured interviews to examine various aspects of tooling usability, e.g., implementation challenges, perceived benefits, and comparison to other signing tools. This approach let us pose consistent questions across all subjects, with the ability to probe about unique aspects of each subject’s circumstances.
To develop our interview instrument, we did not bind our questions to any prior usability framework. Sigstore is a nascent technology and in such contexts an exploratory design is recommended [59, 58]. Only after this exploratory phase did we map our findings onto a formal usability framework, ensuring it reflected practitioners’ lived experiences rather than constraining data collection to preconceived categories. Therefore, to construct the interview protocol, we:
@@ IV-C Instrument Design & Development
    Reviewed grey literature—Sigstore blogs [67] and usage reports [68, 27, 69, 28, 29, 30, 31, 70]——to ground our questions in real-world implementations and updates.
@@
    Used a snowball review of top security–conference works on signing primitives [11, 71] and empirical studies (see section II), to identify gaps and avoid redundant questions.
Next, we refined our interview protocols through a series of practice and pilot interviews[72]. We first conducted two practice interviews with the secondary authors of this work, followed by pilot interviews with the first two interview subjects. For example, we moved the question about each participant’s team’s major product to the demographics section to provide context for their responses. Our full interview instrument consisted of three main topics (B-D). Topics B and C did not fit into the same analysis framework as Topic D. Thus, following the guidance of Voils et al. [73], we analyzed those topics in another work [BLINDED]. [^1] That work is blinded for review. We attest no duplication of material. In this work, we focus on the data for Topic D.
A summary of the questions in our interview protocol is shown in Table I. Topic D consisted of five main question categories. Questions D3–D4 probe current tooling experiences. Questions D1–D2 explore previous tools, and D5 asks about transition factors. See section -B for the full protocol.
TABLE I: Summary of interview protocol. Data from Topics B and C are analyzed in [BLINDED]. We analyze Topic D in this paper. We present our full protocol in section -B
@@ D. Signing tool adoption (5)
IV-D Data Collection
Population: Our study investigates signing tools using a case study of the Sigstore ecosystem. As such, our target participant pool is centered on Sigstore users. Furthermore, our research questions, as reflected in our interview instrument, aim to understand the reasons prior to adoption that teams choose software tools (Sigstore), how Sigstore is used in signing, and the challenges and benefits of Sigstore.
@@ IV-D Data Collection
To recruit this target population, we employed a non-probability purposive and snowball sampling approach  [74]. We leveraged the personal network of one of the authors to send an initial invitation to members of the 2023 KubeCon (hosted by the Linux Foundation) organizing committee, who were also part of the In-toto [75, 76] Steering Committee (ITSC). This initial invitation yielded 6 participants. We then expanded our recruitment through snowball sampling, relying on recommendations from initial participants (5) and authors’ contacts (7) to recruit an additional 12 participants. We interviewed 18 subjects, but only 17 participants could share enough details for analysis.
Participant Demographics: Our subjects were experienced security practitioners responsible for initiating or implementing their organization’s security controls or compliance. This gives them the relevant context to assess their organization’s strategies for adopting software signing tools. Subjects came from 13 distinct organizations, all companies. Relevant demographics are in Table II. Thirteen participants reported their organizations use Sigstore. Two use internally developed tools and two rely on Notary v1 and PGP signing. Those four participants enrich our findings, as their experiences shed light on usability factors that hinder adoption.
TABLE II: Demographics. For Role, “Senior management” are senior managers/directors/executives; “Technical leader” are senior/lead/partner/principal engineers; and “Engineer” and “Manager” are junior staff.
ID Role Experience Software Type
P1 Research leader 5 years Internal POC software
P2 Senior mgmt. 15 years SAAS security tool
P3 Senior mgmt. 13 years SAAS security tool
P4 Technical leader 20 years Open-source tooling
P5 Engineer 2 years Internal security tooling
P6 Technical leader 27 years Internal security tooling
P7 Manager 6 years Security tooling
P8 Technical leader 8 years Internal security tooling
P9 Engineer 2.5 years SAAS security
P10 Engineer 13 years SAAS security
P11 Technical leader 16 years Firmware
P12 Technical leader 4 years Consultancy
S13 Senior mgmt. 16 years Internal security tool
P14 Research leader 13 years POC security software
P15 Senior mgmt. 15 years Internal security tooling
P16 Senior mgmt. 15 years SAAS
P17 Manager 11 years Security tooling
[…]
IV-E Data Analysis
@@
After transcription and anonymization, our analysis began with a data familiarization stage [77] and proceeded through four phases, carried out by two analysts. [^2] One of the analysts is the lead author. In the first two phases, we inductively memoed and coded our transcripts, developing our initial codebooks without referencing our usability framework [78]. In the final phase, we derived themes from our inductively generated codes and then applied our usability framework to further refine the extracted themes. We describe these steps in detail next. We remind the reader that the transcript memoing and codebook was developed across the full interview transcripts, including Topics B and C. This manuscript focuses on the data and codes pertinent to Topic D.
IV-E1 Memoing & Codebook Creation
@@
First, beginning with the initial 18 transcripts, the two analysts independently memoed a subset of six anonymized transcripts across three rounds of review and discussion, following the recommendations and techniques outlined by O’Connor & Joffe [79] and Campbell et al. [80]. This approach ensured the reliability of our analysis process while optimizing data resources. [^3] Although there is no consensus on the ideal proportion of data to use, O’Connor and Joffe suggest that selecting 10–25% is typical [79]. We randomly chose 6 transcripts (33% of data) to develop our initial codebook. After each round, emerging codes were discussed in meetings, resulting in a total of 506 coded memos. We then reviewed the coded memos and finalized an initial codebook with 36 code categories.
@@
To assess the reliability of the coding process, we combined the techniques outlined by Maxam & Davis [81] and Campbell et al. [80]. Specifically, we randomly selected an unanalyzed transcript, which we coded independently. Following Feng et al.’s [82] recommendation, we used the percentage agreement measure to evaluate consistency, as both analysts had contributed to developing the codebook. We obtained an 89% agreement score, suggesting a stable coding scheme [80, 79]. We resolved disagreements through discussion.
IV-E2 Codebook Revision
@@
Following the reliability assessment of our codebook, and given the substantive agreement recorded, one of the analysts proceeded to code the remaining transcripts in two batches of six transcripts per round. In the first batch, six new code categories were introduced, which were subsequently reviewed by the other analyst. The same analytical process was applied to the final six transcripts, leading to the addition of four new code categories, which were subsequently reviewed by the other analyst. Notably, of the 10 codes added during this phase of codebook refinement, only three provided new insights into the data. By mutual agreement between both analysts, these were categorized as minor codes.
@@
To assess the adequacy of our sample size and analysis, we calculated code saturation for each category in our codebook. Following the recommendations of Guest et al. [83], we measured the cumulative number of new codes introduced in each interview. We saw saturation after interview 12, with no new codes after that. The results are presented in Figure 7.
IV-E3 Thematic Analysis & Usability Framework Analysis
Each interview had a median of 15 codes for topic D. Next, following the method proposed by Braun & Clarke [84], we derived initial themes from our code categories. We focused on identifying themes related to Topic D of our protocol (Table I).
We then conducted another round of analysis using a formative usability framework. We selected Cresswell’s framework (Figure 1) based on the nature of our data.
V Results
First, we present a Sankey diagram illustrating the changes in signing tool usage among our participants’ teams (Figure 4). Many participants (6) transitioned from using no tools to adopting Sigstore. Others transitioned from existing tools to Sigstore: 3 from GPG, 2 from Notary, 1 from Skopeo, and 2 from proprietary or internal tools. The final 4 subjects did not transition to Sigstore.
Figure 4: Practitioners’ Changes in Software Signing Tools.
To understand why they did (and did not) change tools, we structure the results by RQ using Cresswell’s framework.
V-A RQ1: Strengths & Weaknesses of Sigstore
The first aim of our research work was to ascertain from the the experiences of our subjects the strengths and weaknesses they have experienced while using Sigstore.
V-A1 Sigstore’s Strength
We present a summarized list of practitioner-reported advantages of using Sigstore in Table III. First, we extracted the benefits our participants experienced while using Sigstore from their responses. These reflections capture their firsthand experiences with Sigstore.
@@ V-A1 Sigstore’s Strength
TABLE III: Summary of Practitioner-Reported Advantages of Using Sigstore. Practitioners cited Sigstore’s technological and macroenvironmental factors as advantages.
@@
Signing Workflow & Verification P1, P2, P7, P9, P14-17
@@
Use of Short-lived Keys & Certificate 3 –P2, P3, P15
@@
Integrability with SLSA build P16
Integrability with several container registries/technologies P12, P14
Integrability with several cloud-native applications P15
@@
Transparency logs increase security P5, P14
Evaluation of signing adoption using logs P9
@@
Free/Open-Source 2 – P7, P17
We grouped their responses into eight categories: ease of use, identity management for signers, compatibility with emerging technologies, use of short-lived keys and certificates, implementation of a transparency log, encouragement of bundling signatures with attestations, free access to the public Sigstore instance, and reliability of the Sigstore service. When mapped to the Cresswell framework, these factors fell under two categories: Technological and Macroenvironmental.
@@
Ease of use: This was the most frequently discussed advantage of Sigstore over other Software signing tools. Participants commonly highlighted the simplicity of following the Sigstore workflow—particularly how key management is fully automated reduce configuration overhead. P7 (security tools, cloud security) states, “ I don’t really know of any other options that provide the same conveniences that Sigstore provides… It really is just one command to sign something and then one more command to verify it, and so much of the work is handled by Sigstore.” Other advantages associated with Ease of Use, as mentioned by practitioners, include the seamless integration of the Sigstore workflow into CI/CD automation using GitHub Actions and the reduced complexity of key management.
Identity management for Signers: Sigstore’s integration of OIDC identity management was a strength per our participants. In traditional software signing, signer identity is often an ad-hoc feature and may be optional for the signing authority. In Sigstore, however, this behavior is integrated with the Fulcio certificate authority as an identity management step before creating a signature. Participants greatly valued the elimination of the key generation and distribution phase, as it reduced the time spent distributing keys to key servers and searching for keys, streamlining the signing process. This aligns with the intuition that identity management is easier than key management for developers [11]. P16 (SAAS, SSC security tool) highlights this point: “I can associate my identity with an OIDC identity as opposed to generat[ing] a key and keep[ing] track of that key…I could say, ‘Oh, this is signed with [Jane]’s public GitHub identity’…that’s much better”
Compatibility with Several New Technologies: A portion of our participants reported that Sigstore’s integration and compatibility with a wide range of modern technologies provided a significant advantage. The mentioned technologies included security tooling such as SLSA build attestations, as well as containerization technologies and cloud computing applications. In most cases, the resources required to set up connections to these technologies were minimal or nonexistent. P12 (consultancy, cloud OSS security) states, “a lot of clients are getting into signing their containers pretty trivial to do if you use something like cosign…Sigstore is the easy choice because it just works on any registry, and integrating it to Kubernetes is pretty easy ”
@@
Presence of a Transparency Log: The ability to audit signing actions was crucial for adopters. Participants valued the transparency log in the Sigstore ecosystem because it provides a tamper-resistant record of all changes to an artifact’s metadata—functionality not offered by other signing systems. This auditability gave users greater confidence in the integrity of signed artifacts. {quoting} P9 (SAAS, SSC security): “From a higher level, we’re going to use a transparency log to determine what was actually signed and what was written into that transparency log.”
Other technological factors noted by participants are the reliability of Sigstore’s public instance.
@@
Free/Open-Source: Sigstore’s open-source nature was also highlighted by participants as a key advantage. They noted that this reduces both setup and maintenance barriers, particularly for users leveraging Sigstore’s public deployments. {quoting} P7 (security tools, cloud security): “I think it’s amazing that Sigstore is free and public”
V-A2 Sigstore’s Weaknessses
We summarize the difficulties reported by participants in using Sigstore in Table IV. Most of these weaknesses mapped to multiple usability factors.
TABLE IV: Reported Difficulties Encountered by Participants in Using Sigstore. We indicate the associated usability factor of each category of weakness using – T-technology, P-social/human, O-organizational, M-macroenvironmental factors.